HintsonConfiguringVAX*SystemsforUNIX† Revisedfor4.2BSD:March15,1983 BobKridle ComputerSystemsSupportGroup U.C.Berkeley kridle@berkeley,ucbvax!kridle SamLeffler ComputerSystemsResearchGroup U.C.Berkeley sam@berkeley,ucbvax!sam ABSTRACT Thisdocumentreflectsourexperiencesandopinionsinconfiguringoverthirty VAXestorunUNIX†overthelastfiveyears. Ourprimeconsiderationsinchoosingequipmentare: •Cost •Performance •Reliability •Maintainabilityandmaintenancecost •Deliverytime •Redundancyofthesystem •Conservationofspace,power,andcoolingresources Weconsidercomponentsindividuallyandthendescribeseveralsystempackages builtfromthesecomponents,emphasizingindependentlysingle-sourcesystems, minimizationofcost,andmaximalexpansioncapability. Copyright◯c1983,BobKridleandSamLeffler.Copyinginwholeforpersonal usebysitesconfiguringUNIXsystemsispermitted.Reproductioninwholeorin partforotherpurposesispermittedonlywiththeexpresswrittenconsentofthe authors.Thispaperisbasedonanearlierpaperofthesamenameauthoredby BobKridleandBillJoy. †UNIXisatrademarkofBellLaboratories. *VAX,VMS,MASSBUS,andUNIBUSaretrademarksofDigitalEquipmentCorporation.

Introduction -2- ◯crjk/sjlMarch,1983 DISCLAIMER Thisdocumentsreflectsourpersonalopinions.WeareresponsibleforsoftwareandhardwaresupportofVAXsystems,andtherecommendationswegivereflectwhatwewoulddo.Wearecareful tonotetheequipmentthatwerecommendbutarenotusing;werecommendnosecond-vendor equipmentthatisnotknowntobeinusesuccessfullyatseveralUNIXsites.Inanycaseyoumay getalemon,nomatterwhatyoubuy.Allwepromiseisthatthisiswhatwebelieve.Letus knowwhatyoufindout. WehavelittlefamiliaritywithVMS.Recommendationsmadehereshouldnotbeconstruedtobe applicabletoanyoperatingsystemotherthanUNIX.WehaveoftenadaptedUNIXtothese devicesinawaythatmaynotbepossiblewithotheroperatingsystems. Finally,notethatwehavenotwrittenthisdocumentsolelytosharetheknowledgewehave acquiredwithyou;wehavewrittenitbecausewedonothavethetimetotalktoeveryonewho needsthisinformation.Pleasedonotcallustoconfirmtheinformationhereortoaskquestions aboutouropinions.Wewouldliketohearofyourexperiences,orlearnofmistakesinthisdocumentorproductsthatweknownothingabout,butdonothavetimetochatabouttheinformationthatisgivenhere.Wedowelcomeelectronicmailsenttoouraddressesasshownonthefirst page. PREFACETO1983REVISION TheVAX/UNIXCommunityhasgrownconsiderablysincethelastrevisionofthispaperin mid-1981.Itissafetosaythat4BSDsystemsrepresentasubstantialportionofalltheVAXes sold.BoththehardwareavailableforbuildinggoodVAX/UNIXsystemsandthesupportservices availabletothosedoingthebuildinghaveincreased.Asignificantportionofthesystemsnow beingbuiltarebeingusedinbusinessandprivateresearchenvironmentsinadditiontothosein thetraditionalacademicUNIXstrongholds.Manyofthesesystemsarebasedonbinarylicenses ratherthantheeducationalsourcelicensesofamiliaratuniversities. WehopethisdocumentwillbeofusetoavarietyofpotentialVAX/UNIXusers.EventhepurchaserwhochoosestobuyacompletesystemfromDECoramixedvendorsystemintegrator shouldbeassistedinmakingabetterchoicebyanincreasedknowledgeoftheintegrationdecisions.Intheareaofhardware,youwillfindthatthesectionsondisks,networkinterfaces,and printershavebeenheavilyreworked.Inadditiontothetraditionalemphasisonhardwareselection,inthisrevisionwetrytoatleastprovidepointerstosomeoftheotherrelevantservices available.Alistofhardwareintegratorsweareawareofwhohavedevelopedexpertiseinthe UNIXareaisincluded.Inaddition,somesuppliersof4BSDsoftwareareincluded. Thisrevisiontakesadifferentpolicyregardingtheinclusionofspecificpricesforhardware:No explictpricesareincluded.Wedoocassionallyincludederivedcostssuchasthecostper Megabyteforaparticularsizedisksystem.Theareonlymeanttobeapproximateandarecalculatedfromcurrentquantityonelistprices. Ithasbeenourexperiencethattheinclusionofpricesmadethereleaseofthispapermuchmore difficultwithoutsignificantlyincreasingitsvalue.Wewereobligatedtopassthepaperbyevery hardwarevendorforpricecheckingandevensothecontentsremainedaccurateforonlyabrief period.Inaddition,regionalpricinganddiscountingpolicysvarywidely.Wehavedecidedthat ratherthandependonusforprices,thereadershouldnarrowtheirrangeofchoicesasmuchas Introduction -3- ◯crjk/sjlMarch,1983 possibleandthendosomehardbargainingwiththeirsuppliers.‘‘Everybodygetsadiscount.’’ OVERVIEW Wefirstdiscusscomponents,listingthealternativeswehavetriedandsometimesafewwehave not,andthendiscusssystempackages.WebuyasubstantialportionofourequipmentfromvendorsotherthanDEC.Thereasonsforchoosingsecondvendorequipmentareusuallysomecombinationofmorecurrenttechnology,lowercostforequivalentequipmentorshorterdeliverytime. Wedonotconsiderdevicesthathaveprovenunreliableorwhoseperformanceweconsiderinadequate.*Inaddition,therearemanydevicesthatwehavenoexperiencewith.Asageneralrule, everynewperipheralhasrequiredanon-trivialamountoflegworktogetuptospeed.Wesug-gestusingonlyperipheralsthathavebeenpreviouslyusedsuccessfullyonthetypeofVAXyouare configuring(780,750or730)ordemandingasubstantial(50-100%)discountforbeingaguinea pig.BeespeciallycarefulofUNIBUS†interfaces.AlmosteverymanufacturerofaUNIBUSwidgetnowincludestheVAXasamachineonwhichhisdevicewillwork.Someofthesedeviceshave stillnotbeenwelltestedinthissituation.Theseoftenwillnotworkwithoutsubstantialmodification. Systembuyerswithoutreadyaccesstoanin−househardwarestaffshouldconsidercarefullythe optionofbuyingasmuchDECequipmentaspossible.Ifyouhavethemoneyandtimerequired todothis,therearesomestrongadvantages.OurDECequipmenthas,ingeneral,provensomewhatmorereliablethantheequivalentalternatevendorequipment.Timefromequipmentdeliverytorunningsystemisalsousuallyshorter.DECfieldserviceinourareaisexcellent.Outside serviceavailablefornon-DECperipheralsisspottyatbest. Forsmallerinstallationsthisoptionshouldbecarefullyconsidered.Itiseasierifyoucancallone partyforallyourproblems,ifyoucanaffordit.AtBerkeley,wearewellpasttheinventorylevel whereselfmaintenancebeginstopayoffevenonallDECsystems,sothisisnotaconsideration. Oneofus(Kridle)managesourlocalhardwaresupportgroup. Unfortunately,thelimitedselectionofconfigurationscurrentlyavailablesometimesmaketheall-DECchoicedifficult.ThisisespeciallytrueofthesmallerconfigurationsasDEC’sbottomend peripheralsarelesssatisfactoryforUNIX.Wesaythisnotjustformonetaryreasons;functionally andaestheticallywewouldprefertohaveneithertheRK07disknortheTS11tapeunitinany systemwehavetodealwith. WerecommendgettingfieldserviceatleastonyourCPUforthefirstyear.Ithaspaidoffforus inthecostofpartsalone.Youcandropthecontractaftertheengineeringchangeshavetapered offandmostoftheinfantfailureshaveoccurred.DECrequiresacertainamountofitsperipheral equipmentonthemachinetoqualifyforfieldservice.Weunderstandthatitiscompanypolicy nottoprovideamaintenancecontractforasystemwithoutaDECmassstorageperipheral.If youintendtopurchaseamaintenancecontract,becertainthatyourlocalfieldserviceiswillingto *Anexceptiontothisruleismadewherewehaveyettofindanysatisfactorydevicesinaparticular category.Intheseinstanceswehaveindicatedourreservationsabouttheexistingchoicesinthehopes thatnewproductswilladdresstheproblemswebelieveareimportant.Thereadershouldrealizethat ifavendor’sequipmenthasbeenmentionedinanegativelightitindicatesweatleastthoughthighly enoughofittoevaluateitseriously.Wearenottryingtodamageanycompany’sreputation,merely insurethatimportantinformationissharedequally.

Introduction -4- ◯crjk/sjlMarch,1983 supportatleasttheDECequipmentyoubuy. BANDWIDTHCONSIDERATIONS EvaluationofthedatatransfercapacitiesbetweenthevariouspartsofVAXsystemsisacomplex taskthatplaysacriticalpartinsystemconfiguration.Unfortunately,thereisatremendous amountofmisinformationavailableonthissubjectandlittleusefulharddata.Wehavemade manymeasurementsandarealwaysintheprocessofmakingmore.Whatwecurrentlyknowfollows. The11/780UNIBUSadapteristhedevicemostfrequentlyshroudedinconfusion.DECdocumentsvariouslygivethebandwidthatbetween1.2MB/secand1.5MB/secwhentransferring throughabuffereddatapath.Wearenotawareofanyspecificationsfortheunbuffereddatapath buthavenotbeenabletouseitwithsomedevicesasslowas50KB/sec.Oneexperimentwe conductedinvolvedexaminingtheUNIBUSprotocollineswithascopewhileconstantlytransferringfromadiskdrive.Weobservedthatwhilethedrivewastransferringatanaveragerateof about1.2MB/sectheUNIBUSwasclosetoonehundredprecentbusy.Thistestwasconducted onanotherwiseidlesystem.NootherdeviceswereactiveontheUNIBUSandlargedisktransfers(cylinders)reducedanyregistersetuptimetoaminimum.Weconcludefromthisthat1.2 MB/secistheabsolutemaximumtransferratepossiblethrougha11/780UNIBUSadapter.Our observationsshowedthatthelargestdelayswhiletransferringdataoccurredwhilethebuffered datapathwasbeingloadedorunloadedfromtheSBI.SincetheUBAiscontrolledbyamicro sequencerthatisalsoinvolvedinotherUBAactivitiessuchasprocessinginterrupts,wesuspect thatonanactiveUBAthisbandwidthmaybesomewhatreduced. Measurementsoftheavailablethroughputtoandfromthe4.2BSDfilesystemindicateasignificantdifferencebetweendisksrunningonthenativeprocessorbus(CMIorSBI)andthoserunning ontheUNIBUS.AveragedataratesareconsistentlylowerondisksresidingontheUNIBUS, evenwhenthecontrollerprovidesafewsectorsofbuffering.Thisleadsustobelievethatwhen averagereadsare4-8Kilobytes(theaverageblocksizeofa4.2BSDfilesystem),mostUNIBUS controllerswillfallbehindandeventuallylosearevolution.Thisdoesnot,however,seemtooccur withtheUDA50UNIBUScontrollerasithasamuchlargeramount(16Kbytes)ofbuffering†. Therearetroublesomedevicesthatcannotbufferenoughdatatoguaranteethatthemaximum sizerecordcanalwaysbetransferred(6250bpitapedrives),ordonotbufferanadequateamount ofdata(RK07diskcontroller).TohandlethesedevicesUNIXprovidesasoftwareinterlockmechanismthatpreventsexcessUBAcontention. TheMASSBUSadaptersarespecifiedtohaveahigherpotentialbandwidthof2.5MB/sec.Since theyareselectorchannelsthatallowonlyonedevicetotransferdataatatime,therealizedbandwidthislimitedtotherateofthefastestdevice.Thefastestdevicescurrentlyavailablefrom DECfor11/750systemsor11/780systemswithasinglememorycontrollertransferat1.3 MB/sec.Large11/780systemswithtwomemorycontrollersandinterleavedmemorymayrun RP07diskdrivesthatthentransferdataat2.2MB/sec.*Aninterestingbandwidthlimitmaybe establishedbythememorycontrollerparticularlyon11/780s.WesuspectthattheCPUmaybe slowedconsiderablybymemorycontentionwhentwodiskchannelsarebeingusedsimultaneously. Thisshouldbealleviatedbyusinginterleavedmemorycontrollers. TheappendixtotheVAXHardwareHandbooktitled‘‘SystemThroughputConsiderations’’seems tobearouttheseimpressionsandshouldbereadcarefullybyanyonehopingtounderstandthe consequencesforVAXapplicationsinvolvinghighbandwidthinputoroutput.Ifwehaddata intensiveapplicationswewouldseriouslyconsidertheuseofRP07disks(andinterleavedmemory controllers)becauseoftheresultanthigherbursttransferrate;thiswillbediscussedfurther †AfewoftheinitialUDA50controllersweredeliveredwithonly4Kbytesofbuffering.Avoidthese. *OnmachineswithonlyonememorycontrollertheRP07hardwareisarrangedtotransferat1.3 MB/sec.

Introduction -5- ◯crjk/sjlMarch,1983 below. MEMORY AllVAXesaresoldwithatleasttheminimumamountofDECmemoryadequatetorundiagnostics.Additionalmemoryisthelowestriskalternatevendorchoice.Webuytheremainderofour 780memoryfromMostek,NationalSemiconductororTrendata.*Thisareaisextremelyprice competitiveandthereareatleastsixpossiblevendors.Byallmeans,askforcompetitivequotes. Assureyourself,however,thatyouarenotthefirstcustomerforanewvendor. AddinmemoryfortheVAX750isaneweritemandpricesarenotaslow.However,thismemoryisalmostidenticaltothe11/70MK11memoryandseveralvendorshavemanagedtobuild thisproductbymodifyingtheirprevious11/70add-inproduct.Trendataalsohas1Mbyte64K RAMmodulesforboth750sand730s. Smallquantities(onetotwomegabytes)areusuallyavailableofftheshelf.Largequantities(4 megabytesandup)havetakenlessthan30days. Forthe11/780memory,theRAMchipsaresocketed,andtworeplacementchipsperboardare suppliedbyallvendorswemention;Youcanpullouttheboardandreplacethechipatyour leisure.Sincesinglebiterrorsarecorrectedthishasneverinvolvedanyunexpecteddowntimefor us.Thereisatleastaoneyearreturntofactoryagreementontheboards,includedinthepurchaseprice.Outofwarrantyrepairsaresaidtotypicallycostlessthan$300.Wehavereturned onlyoneboardtotheplantinabout30boardyears. WhenpurchasedfromDEC,memoryismuchmoreexpensiveforanyofthemachines.Maintenanceona1MegabyteDECmemorymoduleis$179permonthwithboardreplacementthrough fieldservice.Theboardsarenotsocketed.DeliverytimesonmemoryfromDEChavetypically beensubstantiallylongerthantimesfromsecondvendors. Ifyouaregoingtohavemorethan4megabytesofmemoryonyour780youwillneedaCPU expansioncabinetandasecondmemorycontrollerthatincludesasecondhalf-megabyteofDEC memory. Therearetwomodelsof11/750memorycontrollersandbackplanesaround.Theonecurrently beingmanufacturedbyDECcanbefilledwitheitherquarterMegabyteorfullMegabytemodules foramaximumcapacityof8Megabytes‡.Theoldermemorycontrollerandbackplanecanbe populatedwithonlyquarterMegabytemodulesforatotalcapacityof2Megabytes.Tomake mattersevenmorecomplicated,750sexistwhichhavethenewerstylebackplaneandtheolder controller.Thesetoowillonlyhold2Megabytesofmemory.Thesmallercapacitysystemcanbe upgradedtothelargerone,butthisisquiteexpensive;checkwithDECbeforebuyingone,orbe surethatyouwillbesatisfiedwithamaximumof2Megabytes. DISKS† Theareaofdisksanddiskcontrollersisonewhichhasseenagreatdealofchangesincethelast revisionofthispaperinmid1981.AtthattimewehadnoexperiencewithWinchestertechnologydiskdrives.Now,aftersomepainfulexperimentation,wehavesettledonafewWinchester productswhichfillourneedsreliably.Wenolongerbuy,orrecommend,anyremovablemedia diskproducts. *Alistofsecondvendorsandtheirphonenumbersisgivenattheendofthedocument. ‡ItisimportantwhenmixingmemorymodulesizesinVAX11/750stoinstallthememoryinconsecutiveslotsbeginningwiththefirstandstartingwiththe1Mbytemodules. †Disksizesshownthroughoutthisdocumentareinbytesofformattedspaceavailable.

Components -6- ◯crjk/sjlMarch,1983 Thechoiceofavailablecontrollersisalsowiderandmuchimproved.Highqualitycontrollersare availablewhichinterfacetothenativebussesof750sand780saswellastheUNIBUS.Inaddition,DEChasintroducedanentirenewstoragesystemarchitecturewhichplacesagreatdeal morefunctioninthecontroller,incorporatesanewcontroller-driveinterconnect,anduses improvederrorcorrectionalgorithms. First,wewilldiscusssomeofthemajorareasofchangeindisk/controllertechnology.Wewill thenexplorehowtheseimprove,orotherwiseaffect,ourmethodsofdoingbusiness.Finally,we willconsidersomespecificDECandnon-DECproducts. Theavailabilityoflargecapacity,lowcost,highreliabilityWinchestertechnologydiskdriveshas hadanenormousimpactonus.Therackmountable,300Megabyteorbiggerdiskwhichwas always‘‘justaroundthecorner’’isreallyhere.Itishardtoseehowwegotalongwithoutit.We cannowputabout2Gigabytesofstorageinthesamefootprintthatpreviouslyheld256 Megabytes.Inaddition,weconsumeanddissipateabout25%oftheenergywedidwitholder, removablemedia,drives.Theprospectivebuyershouldbewarned,however,thatnotall‘‘winnies’’liveuptoexpectationswithrespecttoreliability.Wearehappywiththereliabilityofthe equipmentwedescribehere.Ifyouwanttotrysomethingelse,besureandhavesomelongheart tohearttalkswithotherusersoftheproduct. CostperMegabyteofdiskstorageisdownsignificantly.Costrangesfrom$30to$110per Megabytefordisks,notcountingthepriceofthecontroller(s).Thisvaluedependsonthesizeof theunitspurchasedandthechoiceofvendor.Costperunitstorageintermsofbothpurchase priceandcosttooperateareastrongerinversefunctionofthetotaldrivecapacitythanever before.Forexample,thecostperMegabyteofthe456MbyteDECRA81isabout35%ofthatof the121MbyteRA80.Thereasonforthisbecomesclearwhenthedrivesareexamined;manyof thecomponentsareidentical. Thehigherrecordingdensitiesofnewdiskdriveshasalsobeenastrongmotivatorincontroller evolution.Onetechniqueforincreasingtherecordingdensityofthedriveshasbeentorelymore heavilyonsophisticatederrorcorrectionandblockremappingschemes.NolargeWinchesterdrive canbedependedontobe‘‘errorfree.’’Infact,mostthedrivesweusehaveuncorrectablemedia defects.Theselocationsmustberemappedusingsomecombinationofcontrollerfirmwareand handlersoftware.Inaddition,thehigherbitratesofnewdiskdrivesdemandfasterseriallogicin thecontrollerinterface.Manyolderdiskcontrollersarelimitedtothebursttransferrateof3330 styledisksofabout1.25Mbyte/sec. Twotypesofcontrollerhaveevolvedforthenewer,highbitdensitydisks.Thefirstissimplya versionofthetraditionalSMDorStorageModuleDriveinterfacereengineeredforhigherdata rates.Thistypeofinterfacecharacterizesallofthenon-DECcontrollerswhichhavebeenpro-ducedforVAXesofthelastfewyears.ThesecontrollersinterfacetothenativebussesoftheVAX (SBIorCMI)wherepossibletoallowthehigherdataratesavailabletobepassedalltheway throughtomemory.Wherethecontrollermustoperateonabusincapableofacontinuoustransferrateashighasthedisk,someamountofinternalbufferingisprovidedtomaximizetheamount ofdatetransferredbeforethedisk‘‘blowsarev’’.* Non-DECcontrollersmostoftenemulatetheDECRH11,RH750,orRH780interface.Somesup-portforerrorcorrectionisprovidedbythecontrolleralthoughasubstantialassistisusually requiredfromthesystemdriver.Remappingofuncorrectablemediadefectsisentirelyhandledby thedriver.All4.2BSDdevicedriverssupportbadblockremapping.Inaddition,errorcorrection andremappingsupportis,optionally,availableinthestandaloneutilities†.Theonlypartofthe **By‘‘blowingarev’’,wemeanadatatransfercannotbecompletedwithoutextraneousdisk revolutions.Thisismainlyafunctionofthetimerequiredbyaprocessortoserviceaninterrupt,the bandwidthofthebus,andthebufferinginthecontroller.Withthe4.2BSDfilesystem,diskcontrollersarenowbeingextendedtotheirlimitations,andbeyond.Thishassignificantlyinfluencedour concernforthetheirlimitationsasbandwidthsuffersgreatlywhensuchaneventtakesplace. †Duetolimitationsinthesizeofabinaryimagewhichmaybeplacedonabootcassetteorfloppy, theerrorcorrectionandbadsectorforwardingcodeisnotincludedinthestandaloneutilitiesby default.

Components -7- ◯crjk/sjlMarch,1983 systemwhichdoesnotgracefullyhandleerrorsormediadefectsisthefirstlevelbootstrapcode usedon750s. DEChasproducedaverydifferenttypeofcontroller,partiallytodealwiththechallengesof higherdensitydiskdrives.Thiscontroller,theUDA50,isanexampleofDEC’slongrangeplan formassstorage(this‘‘plan’’iscalledtheDigitalStorageArchitecture,orDSA).OneofthefundamentalgoalsofDSAistoprovideastandardsetofdisk‘‘operations’’acrossavarietyofstorage products.WithDSAitshouldbepossibletoconstructstandarddriverswhichknowverylittle aboutthecharacteristicsorgeometryoftheactualstoragedevicestheyaredealingwith.Inorder tomeetthisgoal,errorcorrection,badblockforwarding,andeventhemappingoflogicalblocks ontothephysicaldiskarehandledinthecontroller.Requeststothecontrollertypicallyconsistof logicalblockaddressesandcounts,alongwithamemorytransferaddress.Responsesthencontain eitherdataorafailuremessage.Thecontrollerindependentlytakesallpossiblemeasuresto recoverdatabeforereturningfailure. Inadditiontoincreasingthefunctionalityofthecontroller,DSAspecifiesanewcontrollertodrive interface.TheStandardDiskInterface,orSDI,iscapableofhandlingthetransferratesofany drivewhichDECmayproduceintheforeseeablefuture.Thisinterfaceisimplementedusingfour electricallyisolatedradialmini-coaxcablestoeachdiskdriveembeddedinatoughrubber-like umbilical. On750and780systemsweare,orwillbe,buyingeitherlarge(404Mbyte)Fujitsudiskdrives andEmulexSBIorCMIinterfacedcontrollers,orDECUDA50controllerswith(456Megabyte) RA81diskdrives.Thechoicehereisnotclearasthetwopackagesarebothattractiveandeach hasadifferentsetofadvantages.AlthoughwedonotcurrentlyhaveanyUDA50/RA81sat Berkeley,severalusersof4BSDdohavethem,andareverysatisfied.Inaddition,wehavevisited ColoradoSprings,wherethedrivesaremanufactured,andrunbenchmarksonthemusinganearly versionof4.2BSD.Thepreliminarymeasurementssupportouroptimismaboutthe UDA50/RA81combination,thoughwearenotyetreadytopublishtheseresults(theywillbe availableatalatertime). Itisimportantnottoplacetoomuchemphasisonrawperformanceissueswhencomparingproductsassimilarincapabilitiesasthelargediskchoicespresentedhere.Reliability,freedomfrom bugs,andeaseofmaintenanceareequallyifnotmoreimportanttous.Thevalueoftheproduct infutureconfigurationsisalsoimportant.Forexample,theUDA50/RA81disksystemrepresents anearlyimplementationofanewarchitecture.Itincorporatesmanynewfeaturesheretofore unavailabletous.Inaddition,itisexpandableinthesensethatthedisk/controllerinterfaceis designedtohandlefuturedensityincreaseswhicharenotlikelytobeuseablewiththetraditional SMDinterface.Ontheotherhand,anyimplementationasnewastheUDA50/RA81isnotas likelytobeasbugfreeoraswellunderstoodasthetraditionalRHstyleinterfacearchitecture. Table1indicatessomeofthetradeoffsaswenowunderstandthem. Whensearchingforlessstorageforsmallersmallersystems,orwheretwoarmsareneededforperformanceand800+Megabytesofstorageisoverkill,anotherchoiceisrequired.Evenat $50/Mbyte,a404Megabytedriveisnotcheap.Oneoftheauthorshashadgoodexperienceona small750systemwitha160MbyteWinchesterdiskdrivefromTecstoreandaNationalSemiconductorHEX-3000combinationtapeanddiskcontroller.Wealsoknowofsuccessfuluseofthe SpectraLogiccombinationcontrollerona730system.Usingslightlylessexpensivediskdrives andacombinationcontrolleronecanobtaincosteffective(<$75.00/Mbyte)storageinsmaller amountsandprovideatapeinterfacetoboot(sotospeak.) TAPES WeuseEmulexTC-11/PUNIBUStapecontrollersandKennedymodel9300-3800/1600BPI125 IPStransports.CiphertapedrivesandWespercocontrollersarealsowidelyused.Whenpurchasingsecondvendorequipment,onewillalsoneedcablesandarackinwhichtomountthetape Components -8- ◯crjk/sjlMarch,1983 CriterionUDA50/RA81EmulexSC7?0/FujitsuEagle InitialPurchase Cost−750 UDA50and1stRA81− $57.00/Mbytew/oadditional UNIBUSadaptor;$70.00/Mbyte withUNIBUSadaptor SC750andfirstEagle− $55.00/Mbyte InitialPurchase Cost−780 UDA50and1stRA81− $83.00/MbytewithUNIBUS adaptor SC780and1stEagle− $65.00/Mbyte CostforIncrementalAddition AdditionalRA81s− $41.00/Mbyte AdditionalEagles− $32.00/Mbyte Performance Maybesomewhatbetterin mixedrequest,multidriveenvironmentduetoorderingoptimizationspossibleincontroller; softwarehandleratpresentis suboptimal Initialtestsindicate5-10%bet-tersinglefilethroughputdueto bettersustainedburstrate Maintenance Costs Verylow−$111/Mo.for1st driveandcontroller(compareto $326forRM05)

Unknownbutbelievedverylow MeanTime BetweenFailure Toolittleexperienceavailable yet;RM80isprecursorofRA81 mechanicallyandhasbeenquite good Notalotofexperienceonthese yeteither;initialexperience looksexcellent(smallerFujisare phenomenal;30,000MTBF!) MeanTimeto Repair Designedforquickfieldremoval ofHDA;easytorepair Notaseasy;morecomplexdisassembly Sourcesof Maintenance DEC;maint.contractcheap, real,andavailable Notsoclear;askforexchange contractfromvendor Robustnessof DriveInterconnect Incredible−electricalisolation andyoucouldrunovercables withaforklift!Radialconnectionallowseasyremovalofa singledrive SameoldSMDflatcables;daisy chain FutureValue Earlyimplementationofnew architecture;ifitpansout,likely tobecompatiblewithfuture, highperformance,products; DECresalehighanyway Highperformance(stretchedto limits)implementationofold interfacestandard;notlikelyto workagainfornextincrease CosttoIntegrate Handlerisnew;someinitial bugslikely;probablyabugor twoleftincontrollerfirmware too Wellknowninterface;much morelikelytobebugfree Table1.LargeDiskSystemComparison drive.TheKennedytransportcomeswitha15monthfactorywarranty.Ourdistributor exchanges/repairsthecardsinthecontrollersbasedonalocaldiagnosticmodeinthetransport. Afterthewarrantyperiod,cardswapscostabout$75.Fortransportmechanicalfailuresthe transportisreturnedtothefactoryinMonrovia,California,orwefixitourselves.

Components -9- ◯crjk/sjlMarch,1983 GeorgeGobleatPurdueisusinga6250tapesystemwithUNIX.ItincludesaTelex6253drive (800/1600/6250BPI)125IPSwithaTELEXFormatterandanAviv1boardUNIBUSinterface. TheUNIBUSinterfacehas4KBofbuffering,tohelpwithbuslatencyproblems,anditreally appearstobenecessary.Thewholesystemcosthimaboutthreetimeswhatour1600bpisystems cost.TheAvivcontrolleremulatesaTU10whichissimilartotheEmulexNRZ/PEcontroller. Whenheavydatatransferisdonetothedriveat6250bpiitusestheentirebandwidthofthe UBA.ThisforcesUNIBUSaccessthroughtheUBAtobearbitratedbytheoperatingsystemin orderthatthetapedriveandadiskcontrollermaycoexistonthesameUBA.N.B.:Thedriver forthiscontroller/transportcombinationisnotcurrentlyincludedinthestandard4BSDsystem butistriviallyclonedfromtheTM11handlerwhichisastandardpartofthedistribution.Aviv alsohasaTM-11compatiblecontroller,theTFC822,whichsupportsbothKennedyandCipher transports.ThiscontrollerhasmoreinternalbufferingthantheEmulexTM-11emulatorandmay bepreferableforthisreason. NameSpeedDensities Kennedy125ips800/1600 Telex125ips800/1600/6250 OuroriginalVAXsystemcameinapackagewithaDECTE16onitsownMBA.TheTE16is reliablebutslow.TheDECTU45isfaster,butfraughtwithproblemsasthehighmaintenance costreflects.TheDECTU77isagoodtransport,buttheauto-loadingfeaturesdonotseemto workwell,anditisexpensive.Finally,thereisarelativelynewproductfromDEC,a 1600/6250bpi125ipstapedrive,theTU78.ThisisthesametransportastheTU77.Wehavetwo TU78sinuseoncampuswithmixedresults. TheUNIBUStapedrive,theTS11,isincludedinpackagesforthe11/750exceptfortheRK07 packagesystem.Itdoesnothaveavacuumcolumn,andisthushardontapes.Itisaproblemto loadandhasbeenfoundtobeunreliable. NameSpeedDensities TS1145ips1600(Notrecommended) TE1645ips800/1600 TU4575ips800/1600(Notrecommended) TU77125ips800/1600 TU78125ips1600/6250 TERMINALINTERFACES WithaVAXyouget8linesofDZ-11thatprovidesomemodemcontrolbutarenotDMA.We usetheAbleDH-11emulator,theSuperMAXDH/DM,oroneofthetwoEmulexDH-11emula-tors−theCS-11orCS-21.WealsohavetriedtheIntersilDH-11emulatorandknowittofunc-tionsatisfactorily.AlloftheseprovideDMAonoutputandmodemcontrol.TheCS-11is unusualinthatitprovidesexpansionofuptofour16lineDHsonasingleUNIBUShexmodule byplacingtheRS-232supportandUARTSoutonthedistributionpanelsandbussingthesepan-elstotheUNIBUSmodulewithoneribboncable.TheCS-11isanattractivesolutionwherea verylargenumberoflineswillbeconnectedtoonemachinesinceitreducesthenumberofcables, andUNIBUSbackplanespaceandpowerrequired. 4BSDalsoprovidesupportfortheasynchronousserialportionofthetheDECDMF-32.Thisis thestandardcommunicationsinterfacefortheVAX11/730andhasanadditionalfeatureofsupportingbothDMAandprogrammedinterruptoperationforbothinputandoutput.The4BSD drivercurrentlydoesnotuseallthisflexibility,treatingitprettymuchlikeaDH-11.The DMF-32driveralsoworkswiththeAbleDMZ-11,aproductwhichemulatestheasynchronous Components -10- ◯crjk/sjlMarch,1983 serialportionsoftwoDMF-32s. Intheareaofnon-DMAcontrollersfromDEC,therearetheDZ-11andDZ-32(aDZ-11withfull modemcontrol). BoththeDZ’sandtheDH’shaveinputsilo’sthatUNIXcanusetoreduceinterruptloadon input.TheDMAoutputoftheDHemulatorsisespeciallyimportantforgraphicsapplications wherehigh-volumeandcontinuousoutputoccurs. PRINTERS Oneofthemostexcitingdevelopmentsintheareaofprintersistheavailabilityofdesktoplaser printers.ThispaperwasprintedonanImagenlaserprinterwehavebeenusing,quitesuccessfully,forseveralmonthsnow.TheImagenoffershighresolution(240dots/inch),usesplainpaper, andseemstorequireminimalhardwaremaintenance.ItisinterfacedtooneofourVAXesviaa 19.2KbaudRS-232linealthoughaparallelinterfaceisalsoavailable. AmongtheproblemswiththeImagenarethesmallnumberofavailablefontsandtheincompletenessofsomeofthosewhichareavailable.Inaddition,theCannonLBP-10printingengineused hasonlya200sheetpapertray.SincetheunitemploysawetprocessXerographyandsmellsa bit,itisnotlocatedinthesameroomasapersonwhomightberesponsibleforrefillingthetray. Thisinevitablyresultsinprintjobsbackingupinalongqueueuntilsomeonenoticespaperis needed.TheImagenfolkswereinitiallyTEXorientedandtheirtroffsupportcontainsglitches whicharepurportedtogoawaywithfuturereleasesofthesoftware.Wealsohopetoeventually interfaceourprinterdirectlytotheEthernet;assoonasImagenprovidesthenecessarysoftware todoso. AnotherlaserprinterbasedontheCanonLBP-10engineisproducedbySymbolics.Symbolics offersbothRS-232andparallelinterfacestotheprinter.TheSymbolicssoftwareisknowntopro-videexcellentsoftwaresupportfortroff.WearearenowevaluatingaSymbolicsprinter. QMSinGeorgiahasapparentlysolvedthemysteriesoftheXerox2700printerandisdistributing anOEMversionwhichmightbeagoodchoice.Themajorpotentialadvantagesherehavetodo withXerox’ssizeandextensivefieldsupport.Theunitisdryprocess(unliketheImagenand Symbolics)andhas300dots/inchresolution.Withanyluck,wewillalsobeevaluatingthisunit soon. WehavebeenusingsomePrintronix300and600lineperminutedot-matrixprinters.ThePrint-ronixprintersdopoint-plottingat60pointsperinch.Theyarenotoutstandinglycheap,butare ruggedlybuilt. ThenewDataProductsB-600-1isa600LPMbandprinter.Wehaveoneandarebuying another.Althoughwehadsomeinitialproblemsgettingthefirstunitintoservice,itnowruns reliablyandisourheaviestusageproductionprinter. PLOTTERS Electrostaticprinter/plottersthatarecapableof200dots/inchareusablebothasplottersandas outputdevicesfortroff.WehaveanoldmodelVarianthatrequiresconsiderablecareandfeeding; newermodelsaresaidtobelessofaheadache.AnewVersatec11’’modelsellsforabout$8,000. Theobjectionstoalltheseguysarethatthepapertendstobewetsometimes,stinky,andmore expensivethanlineprinter($20per1000sheets).Thesearehighmaintenanceitemsasareall heavilyusedhardcopyoutputdeviceswearefamiliarwith.Fortroff,wenowvastlypreferthe Components -11- ◯crjk/sjlMarch,1983 Imagenlaserprintermentionedabove. NETWORKINTERFACES Networkscanbecategorizedaslocalareanetworks(LANs)orlonghaulnetworksaccordingto theirgeographicallimitations.ThemostwidelypublicizedlocalareanetworkistheEthernet.An exampleofalonghaulnetworkistheDARPAInternetwhichspansmanycontinentsandincludes devicessuchascommunicationsatellitesforconnectingdisjointsub-networks. Amonglocalareanetworksthereareseveralcompetingmodulationschemes.TheEthernetand severalothernetworksusesbasebandmodulationtechniques,whilenewertechnologies,suchas broadband,areavailablefromothervendors.Someofthemajordifferencesbetweenbasebandand broadbandtechnologiesaremaximumstationseparation,cablebandwidth,and,currently,perstationconnectioncost.Atthistime,theleastexpensive,andmostreadilyavailablelocalareanetworkinghardwareusebasebandmodulation.However,giventhelimitationsinherentinbaseband modulationschemes,companiesareplacingmoreworkintodevelopinglowcostpartsforusein broadbandnetworks. Asidefromthequestionofbasebandversusbroadband,selectionofmediumisanissue.Coax cableiscommonlyusedbuttypesofcoaxvary.Broadbandnetworksnormallyusethesamestandard75ohmcoaxialcableusedforCATV,whilebasebanduses50ohmcable.Thisimpliesthat upgradinganetworkfrombasebandtobroadbandrequiresexpensiveinstallationofanewcable unlessonethinksahead,oryoursitealreadyhasinstalledcablingforin-houseCATVuse.Fur-ther,thebestmediumintermsofsignallossandnoiseimmunityisfiberopticcable.However, duetoproblemssuchastappingthecable,fewvendorshaveselectedthistechnology.Ifyouplan toconsiderbroadbandatsometimeinthefuture,whileattheoutsetusingbaseband,itiswell worththecostoftheextracabletorunparallel50and75ohmcoax. Inlookingatnetworkcontrollers,wewillconsideronlytheavailablelocalareanetworkinghardware;ourexperiencewithlonghaulnetworksislimitedtotheInternetandsoisofminimalinterest. ThereareatleastfourvendorswithexistingorannouncedEthernetcontrollers,andwiththesoon tobeavailable‘‘Ethernetchips’’morevendorsmayannounceproducts.Itisunlikely,however, thattheEthernetchipswillsignificantlyinfluencethecurrentpricesasthepriceofanEthernet controllerhasalreadybeendrivendownbythemarketcompetition.Whiletheinfluxofnewtechnologymaynotlowercontrollerprices,itissuretoimprovetheirperformanceandreliability. Wecurrentlyuse10Mb/sUNIBUSEthernetcontrollersfrombothInterlanand3Com.Thetwo controllershavealmostidenticalthroughputcharacteristicswith4.2BSD,butneitherhaveproven entirelysatisfactory.The3Comcontrolleristhelessexpensiveofthetwo.Itsdesignisoptimal forsmallPDP-11sandLSI-11swheretheprocessorisresidentonthesamebuswiththecon-troller.Thedesignemploys16or32Kbytesofdual-portedRAMwhichisdirectlyaddressableas UNIBUS(orQ-bus)memory.WhilethisiseffectiveformachinessuchasthePDP-11orLSI-11 wherenopenaltyisrequiredwhenaccessingtheon-boardmemory,withaVAXanymemory accessmustbearbitratedbytheinterveningUNIBUSadaptor.Theresultofthisisthataccesses totheon-boardmemoryareheavilyconstrainedbythecharacteristicsoftheUNIBUSadaptor. InaccessingmemorythroughaUNIBUSadaptor,allaccessesmustbeperformedonevenbyte boundariesandbenomorethantwobytesatatime.Consequently,onemusteitherbeverycarefulaboutthecodingofanetworkinterfacedriver,orthecontentsofanyon-boardmemorymust becopiedintomainmemorybeforemanipulatingit.Duetothearchitectureofthenetworking subsystemincludedin4.2BSDandthelackofcontroloverthecodegeneratedbytheVAXCcompiler,constrainingmemoryfetcheswasinfeasibleandthesecondalternativewastaken.This impliesthatdatamustbeblockcopiedintoandoutoftheon-boardmemoryawordatatime. TheVAXmovc3instructionisnotusableintheUNIBUSaddressspace,makingthisanexpensive operation.

Components -12- ◯crjk/sjlMarch,1983 Asecondproblemwiththe3Comcontrolleristhatitlacksanon-boardtimerforimplementinga backoffalgorithmwhenaccessingtheEthernet.Thisimpliesthehostmustperformatimingloop whenbackingofffromacongestedEthernet.WhenanEthernetisheavilycongestedthismay provetobeverycostlyasnootherprocessingmaytakeplacewhilethehosttimingloopisexecuting. Athirdproblemwiththe3Comcontrolleristhatitdoesnotallowahosttoreceiveitsown broadcastpackets.Thisimpliesthatbroadcastpacketsmustbecapturedinsoftware.Weconsiderthisaseriousdeficiencyasitpreventshardwaretestingwithoutanauxiliaryechoserver. ThesecondEthernetcontrollerwehaveusedismadebyInterlan.ThiscontrollerprovidesDMA access,aswellasseveraldesirablefeaturessuchason-boardretransmissions.Unfortunately,while theDMAinterfaceshouldbeexpectedtoprovidehigherthroughputthanthesharedmemory approach,usingtheInterlaninterfacewehavebeenabletoattainonlycomparabletransferrates tothosemeasuredwiththe3Cominterface.Inaddition,thecontrollerconsumesasignificant amountofof+5voltpower.Whilebroadcastpacketsareretrievedbytheinterface,theEthernet CRCcalculationisnotperformed. WeknowoftwootherEthernetcontrollers,onefromACCandonefromDEC.WehavetwoACC controllersforevaluation,buthaveyettogainanyexperiencewiththem.TheACCcontrolleris basedontheUMC-Z80andprovidesaDMAhostinterface.TheDECEthernetcontrollerwas announcedatthelastDECUSmeeting,butasofyetweknowofnoneincustomerhands. TosummarizetheEthernetcontrollersituation,itappearsthebeststrategytofollowistowait forEthernetchipstobecomewidelyavailablesothevendorscanreengineertheirexistingcontrollerswithminimalcost.IfyourequireEthernetaccessfromyourVAXnow,youmaywishto followourapproach:selectthelowestpricedproductandtreatitas‘‘disposable’’intheexpectationthatsomethingbetterwilleventuallybeavailable. OtherthanEthernet,theProteonproNET10Mb/sringnetworkisalsopopular.Thisdeviceis alsoknownastheVersionIIlniringnetworkandisinheavyuseatLBLandMITwithgood results.TheProteonproNEToutperformsboththe3ComandInterlancontrollersmentioned aboveinthroughputbenchmarksrunwiththe4.2BSDnetworkingsupport.Further,thering designeliminatesthestandardcomplaintsaboutringarchitecturesbyuseofastar-shapedring configuration.Thestar-shapedringallowseasyadditionanddeletionofnodeswithoutsplicing drillingortaping.Also,anynodecanfailwithoutbringingdowntheringbecauseitisbypassedat thestar-shapedring’spassivewirecenter.Themajorconcernwitharingnetworkisthatitis incompatiblewiththedefactostandardEthernet.Costperstationisslightlyhigherthanthe Ethernet,butstartupcostsarelower(unlessyouuseafiberopticwirecenter).Proteonhas announcedtheyareworkingonan80Mb/scontrollerwhichshouldmakethenetworkevenmore attractive. SOFTWARESUPPORT Therehasbeenincreasingdemandfor4BSDatcommercialinstallationsinaformlessexpensive andmoredigestiblethanasourcelicensefromWesternElectricandanunsupporteddistribution fromBerkeley.Anumberofcompanies,licensedbyWesternElectrictosellandsupportUNIXin binaryform,arenowdistributing4BSD.Someofthesecompaniessupport4BSDasanenhancementfortheirhardwareofferingsothersdealonlyinsoftware.LicensesfromthesevendorsnormallycostmuchlessthanaUNIXsourcecodelicense.Thesecompaniesusuallytrytomake 4BSDmorepalatabletothenon-academiccommunitybyprovidingmorefirst-timeuserdocumen-tationandspecializedconsultingaddressingspecificcustomerapplications.Moreformalsoftware supportarrangementsthanthoseofferedbyU.C.Berkeleyarealsoavailable.4BSDsoftware salesandsupportvendorsareincludedinthelistattheendofthispaper.

Packages -13- ◯crjk/sjlMarch,1983 SYSTEMPACKAGES Wenowpresentsomesamplesystempackages.EachrepresentsabalancedsystemfortimesharinguseunderUNIX.PeopleoftenaskushowmanyuserscanbesupportedUNIXintheseconfigurations.Intheabsenceofspecificinformationaboutapplicationstoberun,thisisanunanswerablequestion.Theamountofloadpresentedtothesystembydifferentapplicationsvarieswidely. Wementionwitheachsystemthecountofinteractiveuserstypicallysupportedinouruniversity researchenvironment. Wefirstpresentsystemsbasedon11/750sandthensystemsbasedon11/780s.Witheachexamplewesuggestfunctionallysimilarsystemsconfiguredinatleasttwodifferentways:firstwithas muchequipmentaspossiblefromDECandsecondwiththebestequipmentknowntous.Wewill notconsidertheVAX11/730aswebelieveitisnotaviableoptionformosttimesharingenvironments.Ourexperiencewiththe730indicatesithasapproximatelytherawprocessingpowerofa PDP-11/34sizeCPU.Thus,eventhoughitisareasonablechoiceforpeoplelookingforanentry levelVAX,weconsideritmostlyasingleusermachine. Variousmeasurementsofthespeedofthe11/750and11/780indicatethatthe11/750executesat roughly60percentofthespeedofan11/780.Bycomparison,an11/70runsatroughly75percentofthespeedofan11/780usingthesamebenchmarks,whichinvolvenofloatingpoint,no32 bitarithmeticonthe11/70,andnosystemcalls.ForUNIXtimesharingusagewebelievethat the11/750hasbetterperformancethanan11/70.Thisisduemainlytoadditionaltuningand performanceenhancementstotheVAXkernel,andtothelargeraddressspaceoftheVAXarchitecture. Thefirstsystemweconsiderisasmall11/750.Thisisfollowedbyanexpansionofthe11/750 intoalargersystem.WearefondoftheVAX11/750asitprovidesthemostcomputational powerperunitcostofthethreeVAXimplementations. Thesecondbasesystemisasmall11/780.WeshowhowitcanbebuiltfromaDEC RUA81/TU78packagesystem,andhowtobuilditfrommixedvendorequipment.Wethen expanditintwoincrements. Thesmallsystemswesuggeststartwithasinglediskandtapecontrollerandsomememory.For time-sharingapplicationsweconfigureourVAXsystemsallowing256Kbytesofmemoryforthe kernelandroughlyanadditional100kbytesofmemoryperactiveuser.*Memoryischeap,especiallyforthe11/780,sowedon’tskimponit. Withmorethanafewusers,itiscriticalthatmorethanonediskarmbepresentinthesystem. Thusallbutthesmallestsystemsincludemorethanonedisk.Astheactiveusercountrises,havingmorethanonediskcontrollerisalsoagoodidea.Thelargesystempackagesincludetwodisk controllers.Forreallylargeandi/ointensivesystemswerecommendhighbitdensitydiskdrives liketheFujitsuEagleortheRP07drivefromDECastheyprovideahighertransferratethanthe 1.25Mbytestypicaloftheremainingdrives.Usingthistransferrateeffectivelyrequiresrunning withinterleavedmemory. ItisdesirableonallUNIXsystemstohaveatleast100MBofdiskspacesothatallthesourcefor thesystemandallthestandardprogramsmaybekeptonlinewithsomeroomforlocallydevelopedprograms.Theamountofspacerequiredbyuserprogramsvariesperinstallation;wemanagetorunmanyofourinstructional/researchmachinesusingabout300-600megabytesofspace actively,althoughslightlymorethanthiswouldbedesirable. Ourlargeresearchmachinerunswith1Gigabyteofdiskstorage,with2disksonaUNIBUSand2 disksonMASSBUSadapters.Theweakestpointinthissystemisthatithasonlya45ipsTE16 tapedriveforbackups.Foreventhesmallestsystems,45ipswillsoonseemslow.Wetherefore recommendstartingwitha125ips1600bpitapedrive.Asfull2400foottapereelsholdonly *ThesenumbersworkreasonablywellinanenvironmenttypicalofUniversitywork(coursework, paperpreparation,debuggingprograms,developingapplicationsformicrocomputers,etc.)More demandingapplicationscouldrequiresubstantiallymorememoryperuser.

Packages -14- ◯crjk/sjlMarch,1983 30MBat1600bpi,largesystemsshouldconsiderincludingatleastonetapedrivecapableofwriting6250bpitapes. VAX11/750PACKAGES Wewanttoputtogetherasmall11/750systemcapableofsupportingabout8time-sharingUNIX users,andalarger11/750systemforroughly16-24users.Weneedaminimumof100megabytes ofspaceforthesmallsystemandareasonabletapedrive,preferablya125ipsunitsothattape operationscanbedoneinareasonableamountoftime;ifthesystemistoincludeonlynonremovabledisks,weconsiderthefastertapesystemtobeimportant.Forthelargersystem,we wishtoadddiskspacetogivethesystemaminimumof250megabytesofspace,andhavemore thanonediskarm. Smallsystem Small750System DECSystemMixedVendorSystem 11/750fromBrokerorIntegrator with.50MbyteDECMemorybut8 Mbytecapacity.

CPU11/750 Memory1MbyteDEC1MbyteNational/Trendata/Mostek UDA50UnibusControllerEmulexSC750RH750Emulator RA80121MbyteDriveFujitsu134MbyteDrive DiskSystem EmulexorWespercoController CipherorKennedy125ipstape TapeSystemTGE1645ipsTapeSys. ThesmallDECsystemisbasedontheSV-BXGMB-CApackage,andincludesanRL02inaddi-tiontotheRA80.WebasicallyignoretheRL02whichisoflittleusetousandusethepackage becauseitisthecheapestwaytogetstarted.WeaddaTGE16tapesystemasthebestchoice amongamyriadofevils.Itisreallytooslow,butitisreliableandnottooexpensive.DEChas beenpromisingsomebetterlowcosttapeunitssoon. Themixedvendorsystemisasinexpensiveaspossiblewhileretainingupwardexpandability.If thebuilderweresurethatthissystemwasnotgoingtobeexpandedmuchthenasubstantial amountmorecouldbeshavedfromthecostbymakingseveralsubstitutions.ANationalSemiconductororSpectraLogicsUNIBUScombinationdiskandtapecontrollercouldbesubstituted fortheseparateCMIdiskcontrollerandUNIBUStapecontrollershown.Aslower,perhaps45 ips,tapeunitwithbuiltinformattercouldbesubstitutedforthe125ipstapedrive.Anolder CPUwith2Megabytemaximummemorycapacitycouldbeused.TheseareavailableforsubstantiallylessthantheCPUsequippedwiththenewermemorycontrollerandbackplane.Evenwith thesemodifications,anotherdiskandanotherMegabyteofmemorycouldeasilybeaddedtoproducesubstantialperformanceimprovement.Oneadvantageofthemixedvendorsystemasshown isthattheEmulexSC750controllerkeepsthediskdrivesofftheUNIBUS.IfanEthernetcontrollersisaddedtothesystem,theywillnotbecontendingforthebus. Mediumsystem. Toexpandthisbasicsystemtosupportmoreusers,wewouldaddadditionallines,diskstorage andmemory.Tothesmallall-DECsystemwewouldaddanotherRA80,anotherMegabyteof memoryandaDZ-11E.TothemixedvendorsystemwewouldaddanotherFujitsu134Mbyte Packages -15- ◯crjk/sjlMarch,1983 disk,anAbleorEmulexDH-11emulatorandanotherMbyteofmemory: AugmentingtheSmall750toa‘‘Medium’’System DECSystemMixedVendorSystem AdditionalDiskRA80121MbyteDriveFujitsu134MbyteDrive MoreMemory1MbyteDEC1MbyteNational/Trendata/Mostek MoreSerialLinesDZ-11EAble/Emulex‘‘DH’’ Thereare,ofcourse,furtherexpansionpossibilitiesforthe11/750.Thesevarydependingonthe applicationbutcouldincludeafloatingpointaccelerator,morememoryupto8Mbytes,andan additionalUNIBUSadaptorontheDECsystemifotherhighspeeddeviceslikenetworkinterfaces aretobeontheUNIBUSalongwiththeUDA50.

Packages -16- ◯crjk/sjlMarch,1983 VAX11/780PACKAGES Forasystemwithmoregrowthpossibilitiesthanan11/750,fasterprocessing,andhigheri/o bandwidth,werecommendstartingwithasmall11/780.Ourgoalhereistostartwithasystem capableofsupporting8-16timesharingusersandexpandingthesystemtobecapableofsupport-ingroughly24users.Wealsoconsideralargeexpansionofthissystem,toasystemthatmight support32to40terminaluserstotheexhaustionofavailableCPUcycles.* Smallsystem Foroursmallsystemweuse400Megabytesofdiskstorageanda125ips6250bpitapedrivethat willbecapableofhandlingfilebackupsifthesystemiseventuallyexpanded.Inourfirstexpansionofthissmallsystem,wewishtoaddtotheavailablespacetoaminimumof800Megabytesof diskstorage,acquireatleasttwodiskarms,andaddadditionalterminallines.Inalargeexpansionofthissystemweincludemoreterminals,anadditionaldiskcontrollertogetatleasttwoseparatediskchannels,andanadditional800Megabytesofstorageforatotalof1600Megabytes. TobuildasmallsystemfromallDECequipment,wewouldstartwiththeRUA81/TU78based system,theSV-AXECA-CA.Thissystemincludes8terminallines,4Megabytesofmemory,a 456Megabytediskdriveanda125ips6250bpitape.ThesystemisequippedwithtwoUNIBUS adaptorssothattheUDA50doesnotcontendwithotherUNIBUSdevices.Tothiswewouldadd afloatingpointaccelerator. OnthemixedvendorsystemwewouldsubstituteaFujitsuEagle404Mbytediskdriveonan EmulexSC780SBIinterfacedcontrollerandanAviv/Telex6250tapesubsystem. Small780System DECSystemMixedVendorSystem 11/780fromBrokerorIntegrator with.25MbyteDECMemoryand UNIBUSAdaptorIncluded CPU11/780 Memory4MbyteDEC4MbyteNational/Trendata/Mostek UDA50UnibusControllerEmulexSC780RH780Emulator RA81456MbyteDriveFujitsu404MbyteDrive onownUBA DiskSystem TEU78125ips6250ipsAvivController TapeSubsystemTelexDrive/Formatter TapeSystem SerialLinesDZ-11AAble/Emulex‘‘DH’’ OtherDECFloatingPt.Acc.DECFloatingPt.Acc. Mediumsystem Toexpandthisbasicsystemtosupportmoreusersandgetadditionaldiskspace,wewouldadd additionallinesanddiskstorage. *Usingsystemssimilartothelargestshownhere,inanenvironmentconsistingofsmallstudentprogrammingsomesiteshavereportedrunningupto70interactiveusers;CPUcyclesarethecritical resourcewiththismanyusers.

Packages -17- ◯crjk/sjlMarch,1983 AugmentingtheSmall780toa‘‘Medium’’System DECSystemMixedVendorSystem AdditionalDiskRA81456MbyteDriveFujitsu404MbyteDrive MoreSerialLinesDZ-11EAble/Emulex‘‘DH’’ Largesystem Toformasystemwiththeemphasisonhandlingofdata-intensiveapplications,andtoemphasize totalgrowthofthesystem,wewouldaddaseconddiskchannelandinterleavememorytoincrease i/othroughputandreduceaverageCPUmemoryaccessasmuchaspossible.InboththeDEC andmixedvendorsystemsaCPUextensioncabinetwouldberequiredinadditiontoanother DECmemorycontroller.Wewouldfilloutthesecondmemorysystemto4Megabytes. Formorediskthroughput,wewouldaddanREP07-AA504MBdiskdriveonaMASSBUScon-trollertothebasicDECsystem.Thisdiskprovidesaveryhighburstdatathroughputandcould sharetheMASSBUSAdaptoroftheTapeUnitwithonlyminorperformancelosswhilethetape unitwasbeingused. Toaccomplishthesameendswiththemixedvendorsystem,wewouldsimplyaddasecond EmulexSC780diskcontrollerchannelandatleastonemoreFujitsuEagle404Mbytediskdrive. Augmenting‘‘Medium’’780to‘‘ReallyBig’’System DECSystemMixedVendorSystem AdditionalDiskRP07(516Mbyte)onFujitsuEagle(404Mbyte) andChannelMASSBUSwithtapesys.onanotherSC780controller SecondMemory Controller andCabinet DECDEC AdditionalMemoryDECTrendata/NationalMostek MoreSerialLinesDZ-11EAble/Emulex‘‘DH’’ SUGGESTIONSONBUYINGHARDWARE Theareavarietyofwaysinwhichyoucanacquirethesystemswehavesuggestedhere,whether theybeallDECormixedvendor.Yourchoiceofacquisitionmethodsdependsonanumberof factorsincluding: •Howmuchcanyouaffordtopay? •Howlongcanyouwait? •Howmuchriskandresponsibilityareyouwillingtoassumeforintegratingyourownhardwarecomponents? •Whatkindofmaintenanceisavailabletoyou? •Howmuchhelpyouneedinintegrating4BSD? Hereisasimplifiedbreakdownofthepossibilities: 1.BuyasmuchaspossiblefromyourDECmarketingorganization. Althoughthissolution,inourexperience,takesthelongestandcoststhemost,ithasits advantages.DECislikelytoshipyouawelltested,integratedsystem,closetothetime Packages -18- ◯crjk/sjlMarch,1983 initiallypromised.Inmostcasestheywillsupportyouwellthroughanyinitialstart-up problemswiththehardware.Thesystemboughtthiswaywillautomaticallybeaccepted foraDECmaintenancecontract.Ofcourse,thecan’thelpyoumuchwith4BSD(yet). Also,theyarenotlikelytobeveryflexibleaboutadjustingtheirconfigurationtoyour needs. 2.Buyanall-DECsystemfromaanOEMspecializingin4BSD TheseOEMsarearelativelynewphenomenon.Theyusuallygetamuchbetterdiscount fromDEConhardwareandcanpasspartofthisthroughtoyouintermsofUNIXexpertise aswellasreducedcost.SometimestheywillbeabletodeliverhardwarequicklywhenDEC istellingyoumonths.SincetheyselllargelyDECsystems,youcanstilltakeadvantageof DECFieldServiceandmostsystemssoldthiswayareguaranteedacceptableforaDEC maintenancecontract. 3.Buyamixedvendorsystemfromasystemsintegrator DEChashadalonglove/haterelationshipwithpeoplewhospecializeinbuildingsystems whichuseDEC’sCPUsandothermanufacturersperipherals.Wethinktheseintegrators servemanyusefulfunctions.First,andforemost,theyoftenbuildacheaperandbettersystem,frequentlyonshortnotice.Second,theykeepDEChonest.Sometimeswefeelthey shouldchargefortheirquotations,sincetheseareoftenusedadvantageouslytoencourage DECtocomedowntoamorereasonablepriceonasystem. Don’tassumemixedvendorsystemsarenotmaintainable.Thereisawholespectrumof maintenancepossibilitiesforthesesystems,particularlyinmajormetropolitanareas.Ifyou areconsideringthisroute,besureandspendsometimeonthephonewiththecustomersof yourprospectivevendor.Insistonthenamesoflongtermcustomers,andtalkalotabout maintenanceexperience.Thefolkswementiononthelastpageofthispaperareknownto haveexperiencewith4BSD. 4.Integratethemixedvendorsystemyourself Ifyouarequalifiedforthisadventure,thenyouprobablyknowwhoyouare.Wecan’tbegin totellyouallthepitfalls.Startsmall.Buyamostlyintegratedsystemandaddsomething youcanaffordtohavenotworkforawhile,suchasmorememory(almosttooeasy),ora bettertapedrive,ormoreterminalinterfaces.Ifyoureallywanttodothewholething, findingtheCPUisoneoftheharderparts.GetyourselfacopyofComputerHotLine.You canprobablygetacomplimentarycopybycallingthemat(800)247-2244.Thisisthe socialregisterofcomputerbrokersandasubstantialportionisdedicatedtofolkssellingnew andusedDEC.(HotLine,Inc.alsodistributestheFarmMachineryHotLineandseveral otherclassifiedfleamarketvarietypublications.Theycannotbeexpectedtocontrolthe contentofadds.Useatyourownrisk!) Wewouldliketomaketwomoreobservationsaboutbuyingequipment.Ithasbeenourexperiencethattheserviceyouwillreceivefromyoursourceisdirectlyproportionaltotheriskinusing thatsource.Further,theserviceoftenisinverselyproportionaltothesourcessize.Looselytranslated,littleguysworkharder. ManywhohavedealtwithDECsalesreportdisappointingexperiences.Lackofproductknowledgeandinabilitytobendtocustomerneedsaretypicalcomplaints.Thisisnottosaythatthere arenotexcellentDECsalespeople.Thereare.Andyoumustremember,whenyoufinallyclose thatdealwithyourDECsalesperson,itwillbedelivered,eventually. Ontheotherhand,thesystemsintegratorwhobuildsoneortwosystemsamonthtypicallysucceedsorfailsbasedoftheexperiencesofhissmallcustomerbase.Wehaveknownmanyofthese folkstomakesuperhumaneffortstopulltogetheracustomersystem,oftensucceedingwithout halftheresourcesavailabletoDECsalespeople.Theyarealsomuchquickertopickuptrends likeaninterestin4BSDandstarttomoldtheirservicestofit.Onceagain,thereisalwaysthe exception,the‘‘UnixSystemsIntegrator’’whocouldn’ttellaninodefromatreetoad.Ifyougo thisroute,youhaveagoodselectiontochoosefrom.Spendtimetalkingtopreviouscus-Packages -19- ◯crjk/sjlMarch,1983 tomers. CONCLUSIONS WehavepresentedsampleVAXsystemsoverawideperformancerangeusingbothall-DECand thebestavailablesecondvendorequipment,emphasizing,independently,minimalcostandmaximalexpandability.Usethisinformationwisely;priceshouldn’talwaysbethebottomline. Considertheall-DECsystemifyoucanaffordit.Ifnot,thesecond-vendorequipmentinthe packageshereisallthoughttoworkwellonVAXhardware.Youcanreliablybuildandoperate suchasystem.Whenyouhavestruggledthroughyourparticularsetofdifficultiesandareupand ontheuucpnetwork,besureandwriteusaboutyourexperiences.Goodluck! ACKNOWLEDGEMENTS Thisdocumentrepresentsalotofwork.Itwouldhavebeeneasier,exceptforeveryonewhosent ushelpfulhintsandsuggestionsand,ingeneral,keptushonest.Inparticular,wewouldliketo acknowledgeallthosevendorswhowerepatientwithus,especiallythosewhoseproductswere ultimatelynotincluded.GeorgeGobleatPurduemadeseveralhelpfulcommentswhichgreatly improvedthecontentofthedocument,andhisexperienceswithFujitsuEagleshasmadeasignificantimpressiononus.TheDECDSAengineeringteaminColoradoSprings,includingPaulMassiglia,BillGraceandChuckHesswereparticularlygenerouswiththeirtimeandenergies.Peter WeinbergerofBellLaboratoriessharedhisexperienceswiththeUDA50/RA81withus.Kirk McKusickspenttimetravelingtoColoradoSpringstoaidinevaluatingtheDECRA81diskdrive. DavidMosherhasworkeddiligentlyasthepurchasingagentforCSRGandalsocontributedtoour understandingofthesubtlerpointsofdiskmanufacturingandoperation.JimReedsgavethe paperacarefulproofreadingandfoundmanyoversights.

Vendorreferences -20- ◯crjk/sjlMarch,1983 VENDORREFERENCES ManufacturerProductPhoneVendorcontact 3ComEthernetCont.(415)961-96023Com(MikeHallaburka) AbleAsync.Mux(714)979-7030AbleComputer(NormKiefer) AvivTapecontrollers(619)247-6844Aviv(EdHagenbuch) DataProductsPrinters(415)948-8961MQIAssociates(AveryBlake) EmulexControllers(415)820-2933EakinsAssociates(BobSigal) FujitsuDisks(415)969-5109EakinsAssociates(BobSigal) ImagenLaserPrinters(415)960-0714Imagen(BobWallace) InterlanEthernetCont.(714)752-4002Interlan(GarySteadman) IntersilAsync.Mux(408)743-4300Intersil(AlanTruscott) KennedyTapeTransports(408)245-9291ElectronicMarketingSpecialists MtXinu4BSDBinarySales(415)644-0146MtXinu(BobKridle) NationalMemory(800)538-8514National(DonJohnson) NationalDisk/TapeCont.(800)538-8514National(DonRudolph) NMSDisk/TapeSys(415)443-1669NMS(BobCrippen) PrintronixPrinters(408)245-4392GroupIIIElect.(ScottDrzewiecki) ProteonNetworkCont.(617)894-1980Proteon(AlMarshall) SpectralogicsDisk/TapeSys.(415)443-1669Nat.Mem.Sys.(BobCrippen) SymbolicsLaserPrinter(415)494-8081Symbolics(DavidShlager) TecstoreDisks(408)732-2143Tecstore(MelFeintuch) TrendataMemory(714)540-3605Trendata(MilesEfron) VarianPlotters(408)733-2900Varian(TedDowns) VersatecPlotters(415)828-6610Versatec(BruceFihe) SYSTEMINTEGRATION/SUPPORT NamePhoneContactNotes VLSI(415)490-3555JoeVoelkerMixedVendorSystemsandSupport IDS(408)738-3368DickCavanaughSpecializeinAllDECSystems EakinsAssoc.(415)969-4533BobSigalMixedVendorSystemsandSupport IPS(713)776-0071MixedVendorSystems IversonInc.(415)459-5665JonIversonMixedVendorIntegration UNIQ(415)362-0470AllDECSystems